find (f g)(x) and (g f)(x).
f(x)=-2x^2 + 1, g(x)= -2x
(f g)(x)=

Answer by amarjeeth123(574)

f(x)=-2x^2 + 1, g(x)= -2x
fog(x)=-2(-2x)^2+1=-8x^2+1
gof(x)=-2(-2x^2+1)=4x^2-2.